OpenStudy (anonymous):

The number of books in 11 bookshelves in a library is shown below: 23, 41, 27, 25, 31, 21, 21, 45, 56, 37, 58 Use a histogram to organize the data and answer the following question: Which statement best compares the height of the bars of the histogram? It is equal for intervals 50-59 and 20-29. It is equal for intervals 40-49 and 20-29. It is greater for interval 20-29 than for interval 30-39. It is greater for interval 30-39 than for interval 40-49.

OpenStudy (anonymous):

@orion :)

OpenStudy (mokeira):

draw a graph with the info given

OpenStudy (mokeira):

then use the graph to answer

OpenStudy (anonymous):

Would it be C?

OpenStudy (anonymous):

we can re-arrange the data by means of groupings in terms of ranges... and count the number of books that falls within the range... range 20-29 = {21,21,23,25,27} = 5 range 30-39 = {31,37} = 2 range 40-49 = {41,45} = 2 range 50-59 = {56,58} = 2 what can we say now about the interval based on your choices?

OpenStudy (anonymous):

That range 20-29 has a greater range than the other ones so answer choice C would work, im pretty sure

OpenStudy (anonymous):

yes it's correct, it is the only statement that is true....\(\ddot\smile\)

OpenStudy (anonymous):

Thanks again :D